OpenStudy (anonymous):

What are the approximate solutions of 4x2 + 3 = −12x to the nearest hundredth? x ≈ −3.23 and x ≈ 0.23 x ≈ −2.72 and x ≈ −0.28 x ≈ 0.28 and x ≈ 2.72 x ≈ −0.23 and x ≈ 3.23

OpenStudy (campbell_st):

rewrite the equation as 4x^2 + 12x + 3 = 0 use the general quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a} \] in your question a = 4, b = 12 and c = 3 substitute and evaluate

OpenStudy (anonymous):

180 SR 13.41 2 * 4 = 8 13.41 / 8

OpenStudy (anonymous):

1.67

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

\[x = \frac{ -12 \pm \sqrt{144 - 48}}{8}\] and just use the calculator to get an answer